OpenStudy (anonymous):

Can anyone help??? simplify: cos(126°)cos(36°) + sin(126°)sin(36°)

OpenStudy (mokeira):

ans=0

OpenStudy (anonymous):

I need to know the steps too!

OpenStudy (solomonzelman):

cos(a-b)

OpenStudy (mokeira):

Oh...i just used a calcutator but i know you get the same values only that one side is negative and the other positive so it ends up as 0

OpenStudy (ybarrap):

cos(a-b)=cos(a) cos(b)-sin(a) sin(-b) let a=126 and b=36

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

then what??

OpenStudy (anonymous):

ive gotten that far!

OpenStudy (ybarrap):

cos(126-36)=cos(90)